# Echoes from the Dark: Galaxy Catalog Incompleteness in Standard Siren Cosmology

### v1

@@ 4.4 Correlation with astrophysical parameters
The study of correlations between $H_{0}$ and astrophysical hyperparameters across the different configurations provides insights into how strongly cosmological constraints depend on assumptions about the BBH mass distribution. When the catalog term is not informative, a strong anticorrelation typically appears between $H_{0}$ and the mass scale parameter $\mu_{\rm g}$. This happens because higher $H_{0}$ values shift sources to higher $z$ at fixed luminosity distance, requiring smaller source-frame masses to match the observations. This behavior has been observed in real data (e.g., Abbott et al. 2023c; Mali & Essick 2025; Abbott et al. 2025a), in mock analyses when using photometric galaxy catalogs (Borghi et al. 2024), and is clearly evident in this work as the completeness decreases.
@@ 4.5 Mismodeling hosting weights
We investigate potential biases from incorrect host galaxy weighting assumptions by re-analyzing the mass-weighted mock1 catalog under wrong assumptions $\alpha_{M}=0$ and $\alpha_{M}=2$. These assumptions affect three likelihood components: galaxy weights in the catalog term (Eq. 7), the completeness function (Eq. 11), and the background distribution (Eq. 13). Additional details and supporting figures are included in Appendix C.
@@
Our results are in agreement with previous studies of host galaxy weighting systematics. In particular, Hanselman et al. (2025) and Alfradique et al. (2025) also find unbiased $H_{0}$ constraints in the case of a complete catalog with stellar-mass mismomdeled by equal weighting. In this work, we confirm the existence of a regime where host weight mismodeling does not affect $H_{0}$ constraints also at higher level of incompleteness. The robustness of these constraints is likely due to the assumption of very informative catalog terms with spectroscopic redshift uncertainties. However, assumptions on the GW data can also play a role. Dedicated mock data challenges will be crucial to compare analysis frameworks and better assess these effects.
@@ 5 Conclusions
In this work, we study the cosmological potential of standard sirens, focusing on the impact of galaxy catalog incompleteness and host galaxy weighting on $H_{0}$ constraints. Here we summarize the main results of the paper.
@@
    We build our parent sample starting from the MICECATv2 mock galaxy catalog and describe it with a parametric stellar mass distribution. We then generate three realizations of 100 BBH events at $\mathrm{S/N}>25$ detectable by a LIGO-Virgo-KAGRA O5-like network over about one year of observation. Each realization is based on a different host galaxy weighting scheme: unweighted, weighted by stellar mass, and weighted by stellar mass squared. We explore standard siren constraints achievable with spectroscopic catalogs, simulating incompleteness through stellar mass or magnitude cuts.
@@
    We obtain percent-level constraints on $H_{0}$ even at moderate levels of incompleteness. In the best-case scenario of a complete spectroscopic catalog, precision reaches 1.6%, 1.3%, and 0.9% for no weighting, mass weighting, and mass-squared weighting, respectively. Remarkably, achieving a $2\%$ precision is possible even when the completeness of potential hosts, averaged within the GW horizon, reaches $50\%$. Achieving a $1\%$ precision in the assumed configurations is only possible if BBH hosting probability scales with the stellar mass squared.
@@
    We characterize how the $H_{0}$ constraints degrade with decreasing catalog completeness using a sigmoid function. We find that the completeness threshold and the steepness of the sigmoid transition increase for higher mass weighting schemes, demonstrating quantitatively that missing progressively more low-mass host galaxies has negligible impact on the cosmological constraints.
@@
    We assess systematic effects from mismodeling host galaxy weighting. With complete spectroscopic catalogs, incorrect weighting does not introduce significant $H_{0}$ biases. At moderate average completeness ($\approx 20\%$), stronger-than-true mass weighting still yields unbiased results, while weaker weighting introduces $10\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$ biases. For highly incomplete or empty catalogs (spectral siren regime), mismodeling the weighting in either direction introduces significant biases, highlighting the importance of accurately modeling the host population when galaxy catalog information becomes sparse.
@@
These results are based on some simplifying assumptions. First, we assume that galaxy properties are measured with negligible uncertainty. Second, we assume that we know the true underlying galaxy distribution. Third, we simulate catalog incompleteness with simple stellar mass and magnitude cuts, while actual surveys follow more complex selection functions as they might target different galaxy populations and be subject to different selection effects. In future work, we plan to test the impact of these assumptions by applying realistic survey conditions tailored to ongoing and upcoming large-scale surveys. We also plan to assess the impact of photometric redshift measurements on our results, with particular focus on the trade-off between catalog completeness and the precision of redshift measurements. Together, these steps will be crucial to push the cosmological constraints to higher $z$ and study the expansion history $H(z)$.
Finally, our study assumes a duty cycle of 100% at nominal detector sensitivity for all instruments in LVK O5-like configuration, providing a simple benchmark for expected yearly O5 performance. Under more realistic observing conditions, achieving the same precision on $H_{0}$ would require proportionally longer observation times. Despite these limitations, our results highlight the crucial importance of spectroscopic galaxy surveys for standard siren cosmology. Moreover, our framework demonstrates that even moderately incomplete spectroscopic catalogs can yield competitive cosmological constraints, even when hosting probabilities are not properly modeled.
@@
These findings are particularly relevant for ongoing galaxy surveys such as Euclid (Euclid Collaboration et al. 2025), DESI (DESI Collaboration et al. 2024), and the proposed WST mission (Mainieri et al. 2024). With the increasing number of GW detections, it will be crucial to understand the interplay between the GW sources and host galaxy properties. Looking ahead, third-generation GW detectors like the Einstein Telescope (Punturo et al. 2010; Abac et al. 2025) and Cosmic Explorer (Reitze et al. 2019) will be transformative in terms of detections, potentially revealing more detailed structures in the GW population and correlation with host galaxy properties (Chen et al. 2024). At the same time, space-based detectors like LISA (Amaro-Seoane et al. 2023) will allow us to test these correlations at completely different black hole scales. Together, these advancements will be crucial to unlock the full potential of GWs as cosmological probes.
@@ Acknowledgements.
The authors thank Michele Mancarella and Lucia Pozzetti for helpful suggestions and comments. NB and MM acknowledge support from the grant ASI n. 2024-10-HH.0 “Attività scientifiche per la missione Euclid – fase E”. MM acknowledges support from MIUR, PRIN 2022 (grant 2022NY2ZRS 001). MM acknowledge the financial contribution from the grant PRIN-MUR 2022 2022NY2ZRS 001 ”Optimizing the extraction of cosmological information from Large Scale Structure analysis in view of the next large spectroscopic surveys” supported by Next Generation EU. MT acknowledges the funding from the European Union - NextGenerationEU, in the framework of the HPC project – “National Center for HPC, Big Data and Quantum Computing” (PNRR - M4C2 - I1.4 - CN00000013 – CUP J33C22001170001). GC is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We also acknowledge the CINECA award under the ISCRA initiative for the availability of high performance computing resources through ISCRA-C project LIGEA HP10CXW8OV and HP10CD9GBT (PI N. Borghi). Software: CHIMERA (Borghi et al. 2024; Tagliazucchi et al. 2025), GWFAST (Iacovelli et al. 2022b), numpy (Harris et al. 2020), JAX (Bradbury et al. 2018), emcee (Foreman-Mackey et al. 2013), matplotlib (Hunter 2007), datashader (Bednar et al. 2025), pastamarkers (PASTA Collaboration et al. 2024).
@@ Appendix A Mock catalogs generation
$\Phi(M,z)\,\mathrm{d}M=\begin{cases}\left[\Phi_{1}^{\ast}(z)\left(\frac{M}{M^{\ast}}\right)^{\alpha_{1}}\!\!+\Phi_{2}^{\ast}(z)\left(\frac{M}{M^{\ast}}\right)^{\alpha_{2}}\right]\mathrm{e}^{-\frac{M}{M^{\ast}}}\frac{\mathrm{d}M}{M^{\ast}}&\text{if }M\geq M_{\rm lim}(z)\\ 0&\text{otherwise}\end{cases}$ (24)
@@ Appendix B Results table
Table 2 reports the marginalized $H_{0}$ constraints at different completeness configurations for the three host galaxy weighting schemes studied in this work. For each configuration, we provide the average completeness $\langle P_{\rm compl}\rangle$ computed following Eq. (23). We stress that $P_{\rm compl}$ refers to potential BBH host galaxies not to the full galaxy population (see its definition in Eq. 11). We also include the number of actual host galaxies $N_{\rm host}$ remaining in the catalog after applying completeness cuts.
Figure 9: Bias on marginalized $H_{0}$ constraints at varying completeness levels for 100 BBH detections in an LVK O5-like scenario generated from a mass-weighted catalog (mock 1) assuming three host weighting schemes: unweighted (wrong), weighted by stellar mass (correct), and weighted by stellar mass squared (wrong).
@@ Appendix C Impact of wrong host galaxy weighting
Figure 9 shows the bias in marginalized $H_{0}$ constraints when incorrect host weighting assumptions are applied to the mass-weighted mock1 catalog across various completeness levels. Using the correct weighting scheme ($\alpha_{M}=1$), all configurations yield unbiased $H_{0}$ constraints within $1\,\sigma$. The shift in the median value of $H_{0}$ observed in the spectral siren regime is likely due to the specific realization of the GW events sampled, and such a scatter is expected. Therefore, it is also expected that wrong configurations will be skewed toward this value.
For complete spectroscopic catalogs, incorrect weighting introduces systematic errors that are negligible compared to statistical uncertainties (with $|\Delta H_{0}|<1\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$). Surprisingly, the mass-squared weighting ($\alpha_{M}=2$) yields slightly tighter constraints compared to the true weighting. This indicates that for the configurations studied in this work, a stronger down-weighting of low-mass galaxies reduces the effective number of galaxies contributing to the catalog term without introducing substantial bias.
As the average completeness decreases below $\approx 20\%$, using a stronger-than-true mass weighting still yields unbiased results. This is also expected as $\alpha_{M}=1$ and $\alpha_{M}=2$ configurations have similar background distributions (see Fig. 4, lower left panel). In contrast, using $\alpha_{M}=0$ produces biases of about $10\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$. When catalogs become highly incomplete or empty (spectral siren regime), both incorrect weighting assumptions produce biases, exceeding $1\,\sigma$ for $\alpha_{M}=0$ and $2\,\sigma$ for $\alpha_{M}=2$.

### v2

@@ 4.4 Correlation with astrophysical parameters
The study of correlations between $H_{0}$ and astrophysical hyperparameters across the different configurations provides insights into how strongly cosmological constraints depend on assumptions about the BBH mass distribution. When the catalog term is not informative, a strong anticorrelation typically appears between $H_{0}$ and the mass scale parameter $\mu_{\rm g}$. This happens because higher $H_{0}$ values shift sources to higher $z$ at fixed luminosity distance, requiring smaller source-frame masses to match the observations. This behavior has been observed in real data (e.g., Abbott et al. 2023c; Mali & Essick 2025; Abac et al. 2025b), in mock analyses when using photometric galaxy catalogs (Borghi et al. 2024), and is clearly evident in this work as the completeness decreases.
@@ 4.5 Mismodeling hosting weights
We investigate potential biases introduced by incorrect host galaxy weighting assumptions by re-analyzing the GW events from mock1 using incorrect weighting schemes during parameter estimation (i.e., $\alpha_{M}=0$ and $\alpha_{M}=2$), while keeping the parent catalog and all other assumptions fixed. These weighting choices propagate through the likelihood by affecting individual galaxy weights (Eq. 7), the completeness function (Eq. 11), and the background distribution (Eq. 13). Further details and supporting figures are provided in Appendix C.
@@
Our results are in agreement with previous studies of host galaxy weighting systematics. In particular, Hanselman et al. (2025) and Alfradique et al. (2025) also find unbiased $H_{0}$ constraints in the case of a complete catalog with stellar-mass mismodeled by equal weighting. In this work, we confirm the existence of a regime where host weight mismodeling does not affect $H_{0}$ constraints also at higher level of incompleteness. The robustness of these constraints is likely due to the assumption of very informative catalog terms with spectroscopic redshift uncertainties. However, assumptions on the GW data can also play a role. Dedicated mock data challenges will be crucial to compare analysis frameworks and better assess these effects.
@@ 5 Conclusions
In this work, we study the cosmological potential of standard sirens, focusing on the impact of galaxy catalog incompleteness and host galaxy weighting on $H_{0}$ constraints. Below, we summarize the main results of the paper.
@@
    We build our parent sample from the MICECATv2 mock galaxy catalog and describe it with a parametric stellar mass distribution. We then generate three realizations of 100 BBH events at $\mathrm{S/N}>25$ detectable by a LIGO-Virgo-KAGRA O5-like network over about one year of observation. Each realization is based on a different host galaxy weighting scheme: unweighted, weighted by stellar mass, and weighted by stellar mass squared. We explore standard siren constraints achievable with spectroscopic catalogs, simulating incompleteness through stellar mass or magnitude cuts.
@@
    We obtain percent-level constraints on $H_{0}$ even at moderate incompleteness. In the best-case scenario of a complete spectroscopic catalog, precision reaches 1.6%, 1.3%, and 0.9% for no weighting, mass weighting, and mass-squared weighting, respectively. Remarkably, $2\%$ precision is achievable even when the completeness of potential hosts, averaged within the GW horizon, reaches $50\%$. Achieving $1\%$ precision in the assumed configurations is only possible if BBH hosting probability scales with the stellar mass squared.
@@
    We characterize how $H_{0}$ constraints degrade with decreasing catalog completeness using a sigmoid function. We find that the completeness threshold and the steepness of the sigmoid transition increase for higher mass weighting schemes, demonstrating quantitatively that missing progressively more low-mass host galaxies has negligible impact on the cosmological constraints.
@@
    We assess systematic effects from mismodeling host galaxy weighting. With complete spectroscopic catalogs, incorrect weighting does not introduce significant $H_{0}$ biases. At moderate average completeness (${\approx}20\%$), stronger-than-true mass weighting yields unbiased results, while weaker weighting introduces $10\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$ biases. For highly incomplete or empty catalogs (spectral siren regime), mismodeling the weighting in either direction introduces significant biases, highlighting the importance of accurately modeling the host population when galaxy catalog information becomes sparse.
@@
These results are based on some simplifying assumptions. First, we assume that galaxy properties are measured with negligible uncertainty. Second, we assume that we know the true underlying galaxy distribution. Third, we simulate catalog incompleteness with simple stellar mass and magnitude cuts, while actual surveys follow more complex selection functions as they might target different galaxy populations and be subject to different selection effects. In future work, we plan to test the impact of these assumptions by applying realistic survey conditions tailored to ongoing and upcoming large-scale surveys. We also plan to assess the impact of photometric redshift measurements on our results, with particular focus on the trade-off between catalog completeness and the precision of redshift measurements. Together, these steps will be crucial to push the cosmological constraints to higher $z$ and study the expansion history $H(z)$. Finally, our study assumes a duty cycle of 100% at nominal detector sensitivity for all instruments in LVK O5-like configuration, providing a simple benchmark for expected yearly O5 performance. Under more realistic observing conditions, achieving the same precision on $H_{0}$ would require proportionally longer observation times. Despite these limitations, our results highlight the crucial importance of spectroscopic galaxy surveys for standard siren cosmology. Moreover, our framework demonstrates that even moderately incomplete spectroscopic catalogs can yield competitive cosmological constraints, even when hosting probabilities are not properly modeled.
To put our results in context, current constraints from GWTC-4.0 provide ${\approx}20\%$ uncertainties on $H_{0}$ from 141 CBC dark sirens, using a galaxy catalog where only about one quarter of the galaxies have spectroscopic redshifts (Abac et al. 2025b). Our forecast of $1{-}2\%$ precision with 100 O5-like BBH events and a spectroscopic catalog confirms and extends the findings of our previous work (B24), which achieved percent-level constraints using a complete sample of massive hosts with $\log_{10}(M_{\star}/M_{\odot})>10.5$. Here we validate those results across the full galaxy-mass range and present a comprehensive end-to-end analysis of catalog incompleteness and host weighting effects on dark-siren constraints. Our approach employs realistic mock catalogs, marginalization on the astrophysical population parameters, and hierarchical inference consistent with current LVK cosmology analyses, extending earlier explorations (Gray et al. 2020) and going beyond more simplified analytical models used in recent studies (Perna et al. 2025; Hanselman et al. 2025; Alfradique et al. 2025; Cross-Parkin et al. 2025). Our results further demonstrate that competitive $H_{0}$ measurements will be achievable with an O5-like configuration with moderately complete spectroscopic surveys, and that host weight mismodeling does not necessarily introduce significant systematics on $H_{0}$ even at moderate completeness levels.
@@
These findings are particularly relevant for ongoing galaxy surveys such as Euclid (Euclid Collaboration et al. 2025), DESI (DESI Collaboration et al. 2024), and the proposed WST mission (Mainieri et al. 2024). With the increasing number of GW detections, it will be crucial to understand the interplay between the GW sources and host galaxy properties. Looking ahead, third-generation GW detectors such as the Einstein Telescope (Punturo et al. 2010; Abac et al. 2025a) and Cosmic Explorer (Reitze et al. 2019) will be transformative in terms of detections, potentially revealing more detailed structures in the GW population and correlation with host galaxy properties (Chen et al. 2024). At the same time, space-based detectors such as LISA (Amaro-Seoane et al. 2023) will allow us to test these correlations at completely different black hole scales. Together, these advancements will be crucial to unlock the full potential of GWs as cosmological probes.
@@ Acknowledgements.
We thank Michele Mancarella and Lucia Pozzetti for helpful suggestions and comments. NB and MM acknowledge support from the grant ASI n. 2024-10-HH.0 “Attività scientifiche per la missione Euclid – fase E”. MM acknowledges support from MIUR, PRIN 2022 (grant 2022NY2ZRS 001). MM acknowledge the financial contribution from the grant PRIN-MUR 2022 2022NY2ZRS 001 ”Optimizing the extraction of cosmological information from Large Scale Structure analysis in view of the next large spectroscopic surveys” supported by Next Generation EU. MT acknowledges the funding from the European Union - NextGenerationEU, in the framework of the HPC project – “National Center for HPC, Big Data and Quantum Computing” (PNRR - M4C2 - I1.4 - CN00000013 – CUP J33C22001170001). GC is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. We also acknowledge the CINECA award under the ISCRA initiative for the availability of high performance computing resources through ISCRA-C project LIGEA HP10CXW8OV and HP10CD9GBT (PI N. Borghi). Software: CHIMERA (Borghi et al. 2024; Tagliazucchi et al. 2025), GWFAST (Iacovelli et al. 2022b), numpy (Harris et al. 2020), JAX (Bradbury et al. 2018), emcee (Foreman-Mackey et al. 2013), matplotlib (Hunter 2007), datashader (Bednar et al. 2025), pastamarkers (Collaboration et al. 2024).
@@ Appendix A Mock catalogs generation
[…]
@@ Appendix B Results table
Table 2 reports the marginalized $H_{0}$ constraints at different completeness configurations for the three host galaxy weighting schemes studied in this work. For each configuration, we provide the average completeness, $\langle P_{\rm compl}\rangle$, computed following Eq. (23). We stress that $P_{\rm compl}$ refers to potential BBH host galaxies not to the full galaxy population (see its definition in Eq. 11). We also include the number of actual host galaxies $N_{\rm host}$ remaining in the catalog after applying observational selection cuts.
@@ Appendix C Impact of wrong host galaxy weighting
Figure 9: Bias on marginalized $H_{0}$ constraints at varying completeness levels for 100 BBH detections in an LVK O5-like scenario generated from a mass-weighted catalog (mock 1) assuming three host weighting schemes: constant (wrong), linear (fiducial), and squared (wrong) stellar mass weighting.
Figure 9 shows the bias in marginalized $H_{0}$ constraints when incorrect host weighting assumptions are applied during parameter estimation. We test the mass-weighted mock1 as our fiducial case across various completeness cuts. When correctly setting $\alpha_{M}=1$, all configurations yield unbiased $H_{0}$ constraints within $1\,\sigma$. The shift in the median $H_{0}$ in the spectral siren regime is likely due to the specific realization of GW events. Therefore, wrong configurations are also expected to be skewed toward this value.
For complete spectroscopic catalogs, incorrect weighting introduces systematic errors that are negligible compared to statistical uncertainties (with $|\Delta H_{0}|<1\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$). Interestingly, the mass-squared weighting ($\alpha_{M}=2$) yields slightly tighter constraints compared to the true weighting. This indicates that for the configurations studied in this work, a stronger down-weighting of low-mass galaxies reduces the effective number of galaxies contributing to the catalog term without introducing substantial bias. As average completeness decreases below ${\approx}20\%$, using a stronger-than-true mass weighting still yields unbiased results. This is also expected as $\alpha_{M}=1$ and $\alpha_{M}=2$ configurations have similar background distributions (see Fig. 4, lower left panel). In contrast, using $\alpha_{M}=0$ produces biases of about $10\penalty 10000\ \mathrm{km\,s^{-1}\,Mpc^{-1}}$. When catalogs become highly incomplete or empty (spectral siren regime), both incorrect weighting assumptions produce biases, exceeding $1\,\sigma$ for $\alpha_{M}=0$ and $2\,\sigma$ for $\alpha_{M}=2$.